A mass-spring system undergoes simple harmonic motion on a frictionless surface with amplitude 1.00 (meters) and angular frequency ω=

ω= 6.8 (rad/s).

Calculate the speed of the mass at the point where 1/6 of the total energy is kinetic energy.


1
Expert's answer
2021-07-12T16:27:53-0400

Answer

Total energy

"E=\frac{1}{2}m\omega^2A^2"


And kinetic energy

"KE=\frac{1}{12}m\omega ^2A^2\\\frac{1}{2}mv^2=\frac{1}{12}m\omega ^2A^2\\"


"v=\frac{\omega A}{\sqrt{6}}\\v=\frac{6.8×1}{\sqrt{6}}=2.77m/s"